1 3 +3.2 2 +3 3 +3.4 2 +5 3 +2.6 2 .............................................................................................................................to n terms

13+3.22+33+3.42+53+2.62.............................................................................................................................to n terms

Grade:12th pass

1 Answers

Deepak Kumar Shringi
5 years ago
there are two series here
(2r-1)3and one more is 3(2r)2now do their sigma and apply formula for that
as sigma n =n(n+1)/2
and sigma n^2=n(n+1)(2n+1)/6